OpenStudy (anonymous):

What is the expression that represents 3x^2/11 in radical form?

OpenStudy (anonymous):

\[\Large 3x^{2/11}\]

OpenStudy (anonymous):

\[\bf \sqrt[y]{a^x}=a^{\frac{x}{y}} \rightarrow (3x)^{\frac{2}{11}}=\sqrt[11]{( \ )^2}\] Can you fill in the space? @touseii45

OpenStudy (anonymous):

3x goes inside ?

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

i thought the 3 would go outside the radical

OpenStudy (anonymous):

Nope. Look at what I typed up carefully, that is how it works.

OpenStudy (anonymous):

You should also keep in mind that 3x^2/11 and (3x)^2/11 are two different things. The next time you type up a question, make it clear as to what you mean.

OpenStudy (anonymous):

\[^{11}\sqrt{3x^2}\]

OpenStudy (anonymous):

no @touseii45 you were correct before. "3" is outside the exponent is only for the "x", right?

OpenStudy (anonymous):

im not sure these types of questions are hard for me to understand all it says is Which expression represents \[\Large 3x^{2/11}\]in radical form?

OpenStudy (anonymous):

and it gives me choices

OpenStudy (anonymous):

and only two of them actually make sense since the other two dont have the number 2 in it

OpenStudy (anonymous):

when they meant \(\Large 3x^{2/11}\), then they mean the exponent is only for the "x" not the 3

OpenStudy (anonymous):

okay. is it always like that? unless it has ( ) right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay thanks kid! (:

OpenStudy (anonymous):

so, conclusion, 3 is outside the radical, rest, like what geniud said

OpenStudy (anonymous):

yw